Assume that one year ago you bought 90 shares of a mutual fund for $18 per share, you received a capital gain distribution of $0.38 per share during the past 12 months, and the market value of the fund is now $23. Calculate the total return for this investment if you were to sell it now. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 2 decimal places.)

Total Dollar Return:
Answer #1

Given,

Number of shares bought a year ago: 90

Purchase price per share (P0) = $18

Capital gain received per share (D)= $0.38

Present Market Value (P1) = $23

Therefore, return per share= P1-P0+D = $23=$18+$0.38 = $5.38

Total Dollar Return of the investment= Return per share* No. of shares= $5.38*90 = $484.20